OpenStudy (anonymous):

a)How can you prove that ;Where an object has a steady speed, the slope of the distance–time graph is constant, and the object’s average speed and its instantaneous speed are the same b)What does this statement mean;The three laws of motion allow us to calculate accurately the motion of any object over time and in three dimensions.

OpenStudy (turingtest):

if you have a distance-time graph x(t), what does the derivative x'(t) represent?

OpenStudy (anonymous):

It represents instaneous speed

OpenStudy (anonymous):

at a particular point

OpenStudy (turingtest):

right, and if it says that speed is constant, then \(x'(t)=v(t)=C\) where \(C\) is some constant therefor we can get some function for \(x(t)\) by integrating \[x(t)=\int v(t)dt=\int C dt\]and what do you get by doing that integral?

OpenStudy (anonymous):

distance

OpenStudy (turingtest):

you get the distance function x(t), which after doing the integral you find is....?

OpenStudy (maheshmeghwal9):

from a to b average velocity is this: -